OpenStudy (anonymous):

Decide whether each of the following is true or false: n!(n + 1) = (n + 1)!

OpenStudy (anonymous):

True

OpenStudy (anonymous):

is there any way this needs to be written out? so I can prove it by repeating the steps to reach the answer of (n + 1)! ?

OpenStudy (anonymous):

Google it

OpenStudy (anonymous):

n!(n + 1) = (n + 1)! Factorial pretty much means multiplied by every number less than it until one (0! = 1 though). So, 6! would be 6 * 5 * 4 * 3 * 2 * 1. Notice how this is just 5! times 6. 5! = 5 * 4 * 3 * 2 * 1 5! * 6 = 6* (5 * 4 * 3 * 2 * 1) If instead of 5 we had n and 6 we had n + 1, then this would still hold. So, we could write: (n+1)! = (n)! * (n+1) The only difference between (n+1)! and n! is that the former has one extra term (by definition).